40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 16 17:02:41 2013, Alex Cole, Configuration, Electronics, AS Table Additions photo_1_(1).JPGphoto_2_(2).JPG
    Reply  Wed Jul 17 11:13:36 2013, Alex Cole, Configuration, Electronics, AS Table Additions photo_(1).JPG
       Reply  Wed Jul 17 16:15:42 2013, Alex Cole, Configuration, Electronics, AS Table Additions photo_1_(3).JPGphoto_2_(3).JPGphoto_3.JPG
Message ID: 8862     Entry time: Wed Jul 17 11:13:36 2013     In reply to: 8859     Reply to this: 8863
Author: Alex Cole 
Type: Configuration 
Category: Electronics 
Subject: AS Table Additions 

[Eric, Alex]

For the RFPD frequency response project, we routed the fiber that will connect our REF DET (on the AS table) to our 1x16 optical splitter (in the OMC_North rack), as pictured. (The new fiber is the main one in the picture, which ends at the right edge near REF DET) Note that we secured the fiber to the table in two places to ensure the fiber would remain immobile and out of other optical paths already in place.

At 2:00 we plan to run fiber from our laser module (in rack 1Y1) to our 1x16 optical splitter (in the OMC_North rack) and measure the power output at one of the splitter's output ports. We plan to keep the output power limited to less than 0.5 mW per optical splitter output.

Attachment 1: photo_(1).JPG  129 kB  | Hide | Hide all
photo_(1).JPG
ELOG V3.1.3-